OpenStudy (anonymous):

hector can walk 14 blocks in 6 minutes, how long will it take hector to walk 48 blocks? formula = minutes/blocks x blocks

OpenStudy (anonymous):

(48/14)*6

OpenStudy (saifoo.khan):

there is a more easy method.

OpenStudy (saifoo.khan):

Ratio method! :D

OpenStudy (anonymous):

so it would be 20 min.?

OpenStudy (anonymous):

14/6 = 48/x x = (48 / (14/6)) x= ... mins to walk 48 blocks

OpenStudy (anonymous):

48 min b.c 1 min per block

OpenStudy (anonymous):

im confused /:

OpenStudy (anonymous):

meeru that is absolutely incorrect GloryRae: round it up to 21 (should be 20.5714.....)

OpenStudy (anonymous):

yeah thats what i got

OpenStudy (anonymous):

blocks / min = blocks / min ratio as saifoo said. pug in what you know and solve for x (16 blks / 6 min) = (48 blks / X min) solve for x

OpenStudy (anonymous):

that's just one way of doing it

OpenStudy (anonymous):

thanks you helped me alot!!